OpenStudy (anonymous):

( 3x^8 ) divided ( x^5 ) =

OpenStudy (ajprincess):

\[\large\frac{ a^b }{ a^c }=a^{b-c}\] Similarly \[\large\frac{ 3x^8 }{ x^5 }=3x^{8-5}\] \[=\large3x^3\] Getting it @HorseCrazyGirlForever?

OpenStudy (anonymous):

Sort of. Thanks for your help @ajprincess!

OpenStudy (ajprincess):

Is it confusing?

OpenStudy (anonymous):

Yeah.

OpenStudy (ajprincess):

Which part of it?

OpenStudy (anonymous):

Pretty much all of it. :$

OpenStudy (ajprincess):

Take \(\large a^b\) and \(\large a^c\). Assume that u want to find \[\large\frac{ a^b }{ a^c}\]. The answer will have the same base that is a. But the exponent of the answer will be the difference of the exponents b and c. that is the exponent of the answer will be b-c. So \[\large\frac{ a^b }{ a^c }=a^{b-c}\] Getting it? or still am I confusing u?

OpenStudy (anonymous):

Kind of still confusing... Can you walk me through this one? Usually when I am walked through things, I understand them better. ( 18 x ^ 13 ) divided ( 3x^7 ) =

OpenStudy (ajprincess):

 This can be easily found by the general formula \[\large\frac{ a^b }{ a^c }=a^{b-c}\] \[\large\frac{ x^3 }{ x^2 }=x^{3-2}=x^1=x\] Getting this?